Show step by step:

Find the inverse for y=(x-3)3+1


1
Expert's answer
2021-03-23T06:20:10-0400

"y=(x-3)^3+1"

Interchange the variables:

"x=(y-3)^3+1"

Solve for "y" :

"x=(y-3)^3+1"

"-(y-3)^3=-x+1"

"(y-3)^3=x-1"

"y-3=\sqrt[3]{x-1}"

"y=\sqrt[3]{x-1}+3"

Replace the "y" with "f^{-1}(x)" to show the final answer:

"f^{-1}(x)=\sqrt[3]{x-1}+3"